Let p, q, r and s represent the following: Determine the truth value of each statement.P: 2 is the smallest prime number. Trueq: The capital city of Camiguin is Mambajao. Truer: The sum of 2 odd numbers is odd. Falses: The diagonals of a parallelogram are congruent. False

1
Expert's answer
2021-06-09T14:32:03-0400

P: 2 is the smallest prime number. – True

q: The capital city of Camiguin is Mambajao. – True

r: The sum of 2 odd numbers is odd. – False

let a = 2k + 1 and b = 2m + 1

then a + b = 2k + 1 + 2m + 1 = 2k + 2m + 2 = 2(k + m + 1) which is even

s: The diagonals of a parallelogram are congruent. – False

The diagonals of a parallelogram are not congruent to each other unless the parallelogram is also a square or a rhombus.